Grade 12th passZoology

An earth satellite revolves around the earth in a circular orbit at a height of 300km above the earth surface. If the radius of earth is 6400km, the velocity of satellite will be nearly:
A. 8.1 km/s
B. 7.7 km/s
C. 3.5 km/s
D. 2.4 km/s

Pushpa Raj Joshi
6 Years agoGrade 12th pass

1 Answer

Vikas TU
6 Years ago
Let mass of satellite = m 
Distance b/n satellite and centre of earth , 
=> 6400+300 = 6700 
Centripetal force = Gravitational attraction 
GMm/r^2 = mv^2/r 
GM/r = v^2 
v^2 = sqrt(5.073 *10^10)
=> 7.72860 km/s 
Hope this helps